# Are We Solving a Well-Defined Problem? A Task-Centric Perspective on Recommendation Tasks

### v2

@@ Abstract.
Recommender systems (RecSys) leverage user interaction history to predict and suggest relevant items, shaping user experiences across various domains. While many studies adopt a general problem definition, i.e., to recommend preferred items to users based on past interactions, such abstraction often lacks the domain-specific nuances necessary for practical deployment. However, models are frequently evaluated using datasets from online recommender platforms, which inherently reflect these specificities. In this paper, we analyze RecSys task formulations, emphasizing key components such as input-output structures, temporal dynamics, and candidate item selection. All these factors directly impact offline evaluation. We further examine the complexities of user-item interactions, including decision-making costs, multi-step engagements, and unobservable interactions, which may influence model design and loss functions. Additionally, we explore the balance between task specificity and model generalizability, highlighting how well-defined task formulations serve as the foundation for robust evaluation and effective solution development. By clarifying task definitions and their implications, this work provides a structured perspective on RecSys research. The goal is to help researchers better navigate the field, particularly in understanding specificities of the RecSys tasks and ensuring fair and meaningful evaluations.
@@
copyright: nonejournalyear: 2025doi: XXXXXXX.XXXXXXXconference: Draft; March 2025; ACM Templateisbn: 978-1-4503-XXXX-X/18/06ccs: Information systems Recommender systems
@@ 1. Introduction
In many recommender systems (RecSys) studies, solutions are proposed for a commonly adopted problem: given a set of users, items, and their interactions, the goal is to recommend items that align with users’ interests or preferences. While this general problem definition captures common patterns across recommendation scenarios, its abstraction overlooks critical details needed for practical RecSys applications. Furthermore, discussions on task definition often lack clarity, particularly regarding its scope and practical implications. At the same time, RecSys research is closely tied to real-world applications, where models are primarily evaluated using datasets obtained from operational recommender platforms. The mismatch between abstract problem definitions and domain-specific evaluations leads to inconsistent settings and findings across experiments (Zangerle and Bauer, 2022; McElfresh et al., 2022). In this paper, we present a comprehensive analysis of different RecSys tasks, highlighting their key distinctions and implications for evaluation and real-world deployment.
Recommender systems are not a new research area, and numerous influential works exist. We begin our discussion in Section 2 with a review of several highly cited works in RecSys, emphasizing task formulations. Interestingly, key factors in RecSys were well defined and discussed decades ago. Yet, the community still disagrees on the choice of baselines and datasets (Ivanova et al., 2023; Beel et al., 2024).
One main contribution of this paper is a detailed examination of task definition in Section 3, focusing primarily on the input and output of a mapping function, i.e., a recommender. We discuss a missing element in task formulation i.e., time, and the selection of candidate items for recommendation, both of which have a significant impact on offline evaluation.
Another contribution of this paper is a detailed analysis of the life cycle of user-item interactions in Section 4. In the pre-interaction judgment stage, users are presented with a list of choices or recommended items. Users incur costs when making informed (or sometimes uninformed) decisions to select a recommendation, which serves as a clear indication of user preferences. We then explore the complexity of user-item interactions. For example, some interactions can involve multiple steps, as in e-commerce, where a user may add an item to their cart, proceed to payment, await delivery, interact with the product, and possibly return it. In music streaming, users are not offered a choice but can only skip or fastforward a recommended song. In some cases, true user-item interactions may not be directly observable as in job recommendation.
Based on the task definition and the framework for understanding it, we present our perspectives in Section 5. We consider (i) the ultimate goal of recommendation, i.e., reducing user cost, (ii) the relationship between task formulation, solution, and evaluation, and (iii) the intersection of task specificity and model generalizability. In our discussion, we use a few examples to show the differences in RecSys tasks based on the presented dimensions, i.e., candidate item selection, pre-interaction judgment, and user-item interaction. We hope that this paper provides researchers, particularly those new to recommender systems, with a clearer understanding of the research topic. After establishing a clear understanding, subsequent research can precisely define specific RecSys tasks, select appropriate baselines, and design experiments that account for task specificity.
@@ 2. A Historical Review of Task Formulation
We begin with four of the most widely cited papers, [^1] Citation counts were obtained via a Google Scholar search for papers with the title words “recommender systems” or “recommender system” on February 12, 2025. each with over 8,000 citations. As foundational works in this field, these papers have influenced many researchers, and the tasks they define have likely shaped numerous follow-up studies. The four papers reviewed were published between 2004 and 2011 (Herlocker et al., 2004; Koren et al., 2009; Ricci et al., 2011; Adomavicius and Tuzhilin, 2005).
The Recommendation Problem. In one of the most influential survey papers in RecSys, Adomavicius and Tuzhilin (2005) formally define the recommendation problem as follows:
@@
The authors further note that “the utility of an item is usually represented by a rating, which indicates how a particular user liked a particular item,” a concept commonly known as explicit feedback. Nevertheless, in most subsequent RecSys studies, implicit feedback has become far more prevalent (Rendle et al., 2009). As a result, the utility of an item to a user is often inferred from binary feedback, whether the user has interacted with the item, such as listening to a song, purchasing a watch, or staying at a hotel. Notably, the difference between explicit and implicit feedback primarily affects data modeling and the loss function design when learning $r$, while the core recommendation problem remains unchanged.
Collaborative Filtering. Koren et al. (2009) do not formally define tasks in recommender systems but instead compare the two primary approaches: content filtering and collaborative filtering. They describe content filtering as a method that creates user or item profiles to capture their characteristics, enabling the recommender to associate users with relevant items. In contrast, collaborative filtering relies exclusively on past user-item interactions, such as previous transactions or product ratings. The authors highlight that “a major appeal of collaborative filtering is that it is domain-free, yet it can address data aspects that are often elusive and difficult to profile using content filtering,” further noting that collaborative filtering is “generally more accurate than content-based techniques.”
Due to these advantages, collaborative filtering solutions are often considered more generalizable, aligning well with one of the key objectives in academic research. However, the authors also acknowledge the cold start problem in collaborative filtering, which arises when new users or items have no past interactions (Schein et al., 2002).
@@
Although solutions such as content-based, collaborative, and hybrid filtering are independent of problem definitions, the widespread adoption of collaborative filtering in recent research leads us to assume that user-item interactions remain a key input for typical recommender systems. With this in mind, we reiterate the points made in (Koren et al., 2009), emphasizing that collaborative filtering relies exclusively on past user-item interactions, such as previous transactions or product ratings.
User Feedback and Preference. In the introduction chapter of the RecSys Handbook, Ricci et al. (2011) describe personalized recommendations as ranked lists of items, where recommender systems aim to “predict what the most suitable products or services are, based on the user’s preferences and constraints.” Regarding user preferences, the authors highlight that these can also be inferred from their actions, such as navigating to a specific product page. In this context, user feedback can take multiple forms: explicit feedback, such as ratings; implicit feedback, derived from user-item interactions; and inferred preferences based on observed user behavior.
@@
Task Diversity. Herlocker et al. (2004) provide a thorough discussion on user tasks for recommender systems. Their discussion focuses on end-user tasks (i.e., not marketers or other system stakeholders), [^2] We refer readers to (Zangerle and Bauer, 2022) for more detailed discussion on other users in a recommender system like item provider, platform provider, and other stakeholders. which aligns well with the RecSys tasks to be discussed in this paper. The key user task is to find good items, such as providing users with a ranked list of recommended items. The authors highlight that “there are likely to be many specializations of the tasks within each domain,” and the domain-specific characteristics are reflected in the properties of the datasets, e.g., the need for novelty versus quality, as well as the cost/benefit ratio of false/true positives/negatives. Regarding novelty, in certain domains, “the user goal is dominated by finding recommendations for things a user does not already know about.” In other domains, users expect recommendations to help them find high-quality items among many choices. Regarding cost/benefit, the authors provide an example of the relatively low cost of false-positive movie recommendations. However, in other settings, e.g., if an item comes with a high price tag, then false positives may carry a much higher cost.
Summary. From this historical review, we derive the following main points: (i) The primary task of a recommender system is to suggest items to users based on their preferences. (ii) Collaborative filtering is a widely adopted approach that leverages past user-item interactions. (iii) In most cases, users provide implicit feedback, and their preferences can also be inferred from their actions. (iv) Recommendation tasks can be domain-specific, with various factors to consider, such as the desired level of novelty and quality, as well as the trade-off between cost and benefit.
@@ 3. A Closer Look at the Task Definition
Based on the above discussion and the main points identified, we now examine the RecSys task definition from multiple perspectives. For ease of presentation, we rewrite Definition 2.1 by specifying recommender as a mapping from its input to output.
Figure 1. The classic task definition of recommendation, where user-item interactions are represented in a matrix form. The objective is to predict the missing values, which correspond to unobserved user-item interactions.
@@
Let $U$ be a set of users and $I$ be a set of items. A recommender system aims to produce a ranked list of items for a user $u$, based on user-item interactions $U\times I$.
@@
In this rewritten definition, the recommender system’s input consists of three components: (i) the user $u$ for whom recommendations are to be made, (ii) the set of user-item interactions $U\times I$, which includes all existing interactions made by all users and all items, and (iii) the set of available items $I$ from which recommendations are being made. The system outputs a ranked list of recommended items for $u$, denoted by $R^{u}$.
Following the tradition of illustrating recommender system tasks in matrix form, we present Definition 3.1 in Figure 1. In this figure, the goal is to generate a ranked list of recommended items for $u_{3}$. The items $i_{1}$ to $i_{n}$ collectively form the item set $I$, and the interactions encompass all interactions made between the $m$ users and $n$ items.
Note that this definition is generic to cover all attributes or side information of users or items, because all such information or metadata can be easily derived given a user ID or item ID. As we focus on the core problem definition of RecSys in this paper, we omit all these details.
@@ 3.1. The Missing Input: Time
Recall that in (Koren et al., 2009), the authors note that collaborative filtering relies exclusively on past user-item interactions to make recommendations. In practical scenarios, whether recommending a product for purchase or a song to listen to, it is common sense that recommendations made at a time point $t$ should be based on all information available at or before $t$. Ideally, the user will find the recommendations appealing, then choose to interact with them after the new recommendations are made at $t$.
Adding Time to the Input. With that, we rewrite the mapping function in Definition 3.1 to consider the time dimension. We also make an assumption that each interaction between a user $u$ and item $i$ is associated with a time stamp $t_{x}$ when the interaction occurred, denoted by $(u,i,t_{x})\in U\times I$.
@@
In this rewritten form, we introduce a time point $t$, indicating a recommendation is to be made for user $u$ at time point $t$. We use $(U\times I)_{\leq t}$ to represent all user-item interactions that occurred before time $t$, a simplified form of {$(u,i,t_{x})\in U\times I|t_{x}\leq t$}. Accordingly, the candidate items available for recommendation are those that were registered in the system and accessible to the recommender as at $t$, denoted by $I_{\leq t}$.
@@
Figure 2. The recommendation task along time line. Recommendations are to be made for user $u_{2}$ at time point $t_{e}$. A model is expected to learn from interactions occurred before $t_{e}$ i.e., $(U\times I)_{\leq t_{e}}$ and make recommendation among items available at $t_{e}$, denoted by $I_{\leq t_{e}}$.
While this reformulation may seem trivial and may not impact real-world or online RecSys recommendations, strictly adhering to this task definition poses significant challenges for offline evaluations (Sun, 2023).
Impact to Offline Evaluation. In offline evaluations, user-item interactions in a dataset must be partitioned into training and test sets. The former is used to train the RecSys model, while the latter is used to assess its performance. According to the task formulation in Equation 3, evaluating a model’s performance at time $t$ for a user $u$ requires that the model only access or learn from the interactions that occurred before $t$, i.e., $(U\times I)_{\leq t}$, and generate recommendations from $I_{\leq t}$, the items available at or before $t$.
@@
Figure 2 provides an illustration where the items interacted with by each user are plotted to the right of the user along a global timeline. For example, user $u_{1}$ interacted with item $i_{1}$ at time point $t_{1}$. If we use the user-based leave-last-one-out data split to evaluate our model’s accuracy for user $u_{2}$, the last interaction of $u_{2}$ at time $t_{e}$ is masked as the test instance. The recommender can learn from all user-item interactions that occurred before $t_{e}$, i.e., $(U\times I)_{\leq t_{e}}$. The items available for recommendation to $u_{2}$ at $t_{e}$ are those that have been interacted with by any user before $t_{e}$. Notably, items $i_{7}$ and $i_{8}$ received their first interaction after $t_{e}$, the system has no interaction data for these items at $t_{e}$, and they should not be recommended to $u_{2}$. [^3] Note that our discussion assumes the recommendation model is based on collaborative filtering, to avoid the possible misunderstanding that items $i_{7}$ and $i_{8}$ could be recommended based on profile matching.
Finally, note that the user-based leave-last-one-out data split results in each user having their own $(U\times I)_{\leq t}$ and $I_{\leq t}$, since not all users have their last interaction at a common time. Without enforcing a global time point, many widely adopted data partitioning schemes, such as random splitting or user-based leave-last-one-out, lead to data leakage (Ji et al., 2023). This occurs when user-item interactions from the future (relative to time $t$ in Equation 3 or $t_{e}$ in Figure 2) are used to train the model and make recommendations.
In other words, when recommending items for a specific user, the model does not rely solely on that user’s interaction history but also leverages interactions from other users who may share similar preferences. The time point constrains all user-item interactions accessible to the system.
@@
The incorporation of factor $t$ into the task formulation also directly impacts the implementation of certain baseline models. For instance, the widely used popularity baseline, which simply counts item occurrences in the training data, does not accurately reflect real-life item popularity. In practice, item popularity is typically associated with a reference time point and a specific time range, e.g., the top-selling books from last week or last month. By introducing a time point $t$ in problem definition, item popularity can be derived with reference to that specific time point (Ji et al., 2020, 2022).
@@ 3.2. The Missing Constraints on Candidate Items
Next, we focus on the available items that can be recommended to a user $u$ at time $t$. Herlocker et al. (2004) discuss recommender system tasks from the perspective of novelty and quality, assuming that users generally expect recommended items to be new and previously unconsumed. However, this assumption may not always hold, as a recommender system is ultimately designed to help users find preferred items more easily and quickly.
In many scenarios, users may prefer to interact with items they are already familiar with and confident in (Anderson et al., 2014). Groceries shopping (Ariannezhad et al., 2022; Katz et al., 2022), e-commerce (Wang et al., 2019; Quan et al., 2023), food delivery ordering (Li et al., 2024), and song listening (Tsukuda and Goto, 2020; Reiter-Haas et al., 2021) are a few examples where a user may prefer to have earlier consumed items to be recommended for easy selection.
Let $I^{u}_{t}$ be the items that $u$ has interacted with in the past at time point $t$. We use $I_{t}^{\bar{u}}$ to denote the remaining items available at $t$ that are new to $u$, then the recommendation in Equation 3 can be divided into two formulations for repeated consumption $R_{t}^{ur}$ and for exploration $R_{t}^{ue}$ recommendations, respectively.
@@
Note that, how to present $R_{t}^{ur}$ and $R_{t}^{ue}$ to users through a web or app interface, e.g., as two separate rankings or as a merged ranking of items from both recommendations, is orthogonal to our discussion in this paper.
Figure 3. The recommendation for repeated consumption for user $u_{2}$ at time $t_{e}$. The search space for repeated consumption is the two items $i_{2}$ and $i_{5}$. To recommend new items, the candidate items will be the remaining items in $I_{\leq t_{e}}\setminus\{i_{2},i_{5}\}$.
Impact to Offline Evaluation. Again, the partitioning of candidate items for recommendation $I$ into $I^{u}$ and $I^{\bar{u}}$ may seem trivial (here, we omit the subscript $t$ for clarity), but this partitioning leads to (i) significant differences in the search space for a recommender, because $|I^{u}|\ll|I^{\bar{u}}|$ and $|I^{\bar{u}}|\approx|I|$, and (ii) subsequently significant differences or impacts on evaluation.
It is important to note that if repeated consumption and exploration are not clearly distinguished, the evaluation results may be significantly biased toward repeated consumption. Because of the much smaller item search space for recommendation and a large portion of, if not all, items being users’ preferred items, the chances of making a good recommendation among the repeated items are much higher than picking an item from a much larger pool of items, especially those the user has no experience with. The recommendation accuracy could be easily dominated by the repeated items if repeated consumption is common like food ordering and groceries shopping. Hence, the separation of evaluations of repeated consumption and exploration better reflects the model’s accuracy; a detailed discussion and comparison is presented in  (Li et al., 2024) for food delivery recommendation. Additionally, due to the much smaller search space, recommending repeated items is significantly more efficient than recommending items for exploration.
Candidate Item Selection. The repeated consumption and exploration is just one example of candidate item selection. The candidate items suitable for recommendation can be directly specified by users. For example, if a user is planning a trip to a city, hotel recommendations would only be meaningful if they are located in that city and available during the user’s travel dates. Although this may seem trivial and makes the recommendation problem more similar to a search problem, the vast majority of recommendation work using the Yelp dataset does not consider users’ geographical constraints, to the best of our knowledge. That is, if a user resides in New York City and primarily visits restaurants within a relatively small region of NYC, recommending a restaurant from a different state would be far less practical for this user. On the other hand, the candidate items suitable for recommendation could also be derived based on user intent mining through search logs and other means. Note that the user intent mining here is different from the line of research on intent-aware recommender systems (Jannach and Zanker, 2024; Shehzad et al., 2025), which will be discussed in Section 5.2. One key difference is that, in our discussion on candidate item selection, the set of eligible items to be recommended is determined before running the recommender system, hence is part of the input to the problem definition and can be specified by users, illustrated in Figure 4.
3.3. Yet Another Rewritten Task Formulation
We can now combine Equations 4 and 5 into a more general formulation by introducing a selection condition on the candidate items, denoted as $s(I_{\leq t})$. Specifically, $s(\cdot)$ can represent selection criteria based on user interaction history e.g., for repeated consumption, or be specified by users with temporal, spatial, or other attribute-based conditions. However, by listing $s(I_{\leq t})$ as part of the input to the recommendation task, the candidate set of items is determined before a model ranks them.

### v3

# A Task-Centric Perspective on Recommendation Systems

@@ Abstract.
Many studies in recommender systems (RecSys) adopt a general problem definition, i.e., to recommend preferred items to users based on past interactions. Such abstraction often lacks the domain-specific nuances necessary for practical deployment. However, models are frequently evaluated using datasets collected from online recommender platforms, which inherently reflect domain or task specificities. In this paper, we analyze RecSys task formulations, emphasizing key components such as input-output structures, temporal dynamics, and candidate item selection. All these factors directly impact offline evaluation. We further examine the complexities of user-item interactions, including decision-making costs, multi-step engagements, and unobservable interactions, which may influence model design. Additionally, we explore the balance between task specificity and model generalizability, highlighting how well-defined task formulations serve as the foundation for robust evaluation and effective solution development. By clarifying task definitions and their implications, this work provides a structured perspective on RecSys research. The goal is to help researchers better navigate the field, particularly in understanding specificities of the RecSys tasks and ensuring fair and meaningful evaluations.
@@
copyright: nonejournalyear: 2025doi: XXXXXXX.XXXXXXXccs: Information systems Recommender systems
@@ 1. Introduction
In many recommender systems (RecSys) studies, solutions are proposed for a commonly adopted problem: given a set of users, items, and their interactions, the goal is to recommend items that align with users’ interests or preferences. While this general problem definition captures common patterns across recommendation scenarios, its abstraction overlooks critical details needed for practical RecSys applications. Furthermore, discussions on task definition often lack clarity, particularly regarding its scope and practical implications. At the same time, RecSys research is closely tied to real-world applications, where models are primarily evaluated using datasets obtained from operational recommender platforms. The mismatch between abstract problem definitions and domain-specific evaluations leads to inconsistent settings and findings across experiments (Zangerle and Bauer, 2022; McElfresh et al., 2022).
We begin with a review of several highly cited works in RecSys, emphasizing task formulations. Interestingly, key factors in RecSys were well defined and discussed decades ago. Yet, the community still disagrees on the choice of baselines and datasets (Ivanova et al., 2023; Beel et al., 2024). We then provide a detailed examination of task definition, focusing primarily on the input and output of a mapping function, i.e., the recommender. We discuss the missing elements in task formulation: time and the selection of candidate items for recommendation, and their impact on offline evaluation. Lastly, we review the life cycle of user-item interactions, with a focus on the cost incurred from recommendations made to the feedback of the user-item interactions. Based on the task definition and proposed framework, we outline key perspectives and actionable directions for future work.
@@ 2. A Historical Review of Task Formulation
We begin with a few widely cited RecSys papers (Herlocker et al., 2004; Koren et al., 2009; Ricci et al., 2011; Adomavicius and Tuzhilin, 2005). As foundational works in this field, these papers have influenced many researchers, and the tasks they define have likely shaped numerous follow-up studies. In an influential survey paper, Adomavicius and Tuzhilin (2005) formally define the recommendation problem as follows:
@@
The authors further note that “the utility of an item is usually represented by a rating, which indicates how a particular user liked a particular item,” a concept commonly known as explicit feedback. In subsequent RecSys studies, implicit feedback has become far more prevalent (Rendle et al., 2009). As a result, the utility of an item to a user is often inferred from binary feedback, whether the user has interacted with the item. Notably, the difference between explicit and implicit feedback primarily affects data modeling and the loss function design when learning $r$, while the core recommendation problem remains unchanged.
@@
Koren et al. (2009) compare the two primary approaches in RecSys: content filtering and collaborative filtering. Content filtering builds user and item profiles based on their characteristics, allowing the system to match users with relevant items. In contrast, collaborative filtering relies exclusively on past user-item interactions. Although solutions such as content-based, collaborative, and hybrid filtering are independent of problem definitions, the widespread adoption of collaborative filtering in recent research leads us to assume that user-item interactions remain a key input for typical recommender systems. Regarding user preferences, Ricci et al. (2011) highlight that these can also be inferred from their actions, such as navigating to a specific product page. In this context, user feedback can take multiple forms: explicit feedback, such as ratings; implicit feedback, derived from user-item interactions; and inferred preferences based on observed user behavior.
@@
Herlocker et al. (2004) provide a thorough discussion on user tasks for recommender systems. Their discussion focuses on end-user tasks (i.e., not marketers or other system stakeholders) [^1] We refer readers to (Zangerle and Bauer, 2022) for more detailed discussion on other users in a recommender system like item provider, platform provider, and other stakeholders., which aligns well with the RecSys tasks to be discussed in this paper. The key user task is to find good items, such as providing users with a ranked list of recommended items. The authors highlight that “there are likely to be many specializations of the tasks within each domain,” and the domain-specific characteristics are reflected in the properties of the datasets.
@@ 3. A Closer Look at the Task Definition
For clarity, we rewrite Definition 2.1 by specifying a recommender as a mapping function.
@@
Let $U$ be a set of users and $I$ be a set of items. A recommender aims to produce a ranked list of items for a user $u$, based on user-item interactions $U\times I$.
@@
In this rewritten form, a recommender’s input consists of three components: (i) the user $u$ for whom recommendations are to be made, (ii) the set of user-item interactions $U\times I$, which includes existing interactions made by users and items, and (iii) the set of available items $I$ from which recommendations are being made. The system outputs a ranked list of recommended items for $u$, denoted by $R^{u}$. This definition is generic to cover all attributes or side information of users or items, because all such information can be easily derived from user IDs and item IDs.
@@ 3.1. The Missing Input: Time
In practical scenarios, whether recommending a product for purchase or a song to listen to, recommendations made at a time point $t$ should be based on all information available at $t$. With that, we rewrite the mapping function in Definition 3.1 to consider the time dimension. We also make an assumption that each interaction between a user $u$ and item $i$ is associated with a time stamp $t_{x}$ when the interaction occurred, denoted by $(u,i,t_{x})\in U\times I$.
@@
Here, we introduce a time point $t$, indicating the time a recommendation is to be made for user $u$. We use $(U\times I)_{\leq t}$ to represent all user-item interactions that occurred before time $t$, a simplified form of $(u,i,t_{x})\in U\times I|t_{x}\leq t$}. The candidate items available for recommendation are those that were registered in the system and accessible as at $t$, denoted by $I_{\leq t}$.
@@
Figure 1. Recommendations are to be made for user $u_{2}$ at time point $t_{e}$. A model is expected to learn from interactions occurred before $t_{e}$ i.e., $(U\times I)_{\leq t_{e}}$ and recommend items available at $t_{e}$, denoted by $I_{\leq t_{e}}$.:
@@
While this reformulation may seem trivial and does not impact real-world or online RecSys implementations, strictly adhering to this task definition poses significant challenges for offline evaluations (Sun, 2023). In offline evaluations, user-item interactions in a dataset are partitioned into training and test sets. The former is used to train a recommendation model, while the latter is used to assess its performance. Figure 1 provides an illustration where the items interacted with by each user are plotted to the right of the user along a global timeline. For example, user $u_{1}$ interacted with item $i_{1}$ at time point $t_{1}$. If we use the user-based leave-last-one-out data split to evaluate our model’s accuracy for user $u_{2}$, the last interaction of $u_{2}$ at time $t_{e}$ is masked as the test instance. The recommender can learn from all user-item interactions that occurred before $t_{e}$, i.e., $(U\times I)_{\leq t_{e}}$. The items available for recommendation to $u_{2}$ at $t_{e}$ are those that have been interacted with by any user before $t_{e}$: $I_{\leq t_{e}}=\{t_{1},t_{2},t_{3},t_{4},t_{5}\}$. Notably, items $i_{7}$ and $i_{8}$ received their first interaction after $t_{e}$, the system has no interaction data for these items at $t_{e}$, and they should not be recommended to $u_{2}$. [^2] Note that our discussion assumes the recommendation model is based on collaborative filtering, to avoid the possible misunderstanding that items $i_{7}$ and $i_{8}$ could be recommended based on profile matching.
Note that the user-based leave-last-one-out data split results in each user having their own $(U\times I)_{\leq t}$ and $I_{\leq t}$, since not all users have their last interaction at the same time. Without enforcing a global timeline or an absolute time point, many commonly used data partitioning schemes, like random splitting or user-based leave-last-one-out, can lead to data leakage. Empirically, we show that the impact of data leakage on recommendation models is unpredictable (Ji et al., 2023). Hence, comparing the performance of recommendation models under data leakage in offline evaluation is both impractical and meaningless. Recently, Le et al. (2025) further show that, for sequential recommenders, eliminating data leakage leads to a 21.7 - 73.4% drop in sampled nDCG@10 compared to the commonly adopted setting that includes data leakage. They also observe that the impact of data leakage on model rankings is unpredictable, consistent with our findings in (Ji et al., 2023).
Indeed, some recent studies have realized the importance of this issue, and adopted splits based on absolute temporal cutoffs. To examine how prevalent this practice is, we reviewed the 49 long papers published at the ACM RecSys 2025 conference (September 2025). We found that 16 papers adopted random splits, entirely ignoring the temporal factor. Among the remaining papers that report experimental settings, 7 used relative temporal splits, which also lead to data leakage.
@@
The incorporation of the temporal factor $t$ into the task formulation also directly affects the implementation of certain baseline models. For example, the widely used popularity baseline (which simply counts item occurrences in the training data) does not accurately reflect real-world item popularity. In practice, item popularity is typically defined with respect to a reference time point and within a specific time window, e.g., the top-selling books from the past week or month. By explicitly introducing a time point $t$ into the problem definition, item popularity can be computed relative to that specific time point. Interestingly, the simple temporally aware popularity model, DecayPop (Ji et al., 2020), has been shown to be the most effective ranking method on the Yambda-5B dataset for the music recommendation ‘like’ scenario, outperforming many more complex models (Ploshkin et al., 2025). This finding calls for a careful re-evaluation of prior studies that overlooked the temporal dimension.
@@ 3.2. The Missing Constraints on Candidate Items
Herlocker et al. (2004) discuss recommendation tasks from the perspective of novelty and quality, assuming that users generally expect recommended items to be new and previously unconsumed. In some scenarios, users may prefer to interact with items they are already familiar with and confident in (Anderson et al., 2014). Groceries shopping (Ariannezhad et al., 2022; Katz et al., 2022), e-commerce (Wang et al., 2019; Quan et al., 2023), food delivery ordering (Li et al., 2024), and song listening (Tsukuda and Goto, 2020; Reiter-Haas et al., 2021) are a few examples where a user may prefer to have earlier consumed items to be recommended for easy selection.
Let $I^{u}_{t}$ be the items that $u$ has interacted with in the past at time point $t$. We use $I_{t}^{\bar{u}}$ to denote the remaining items available at $t$ that are new to $u$. Equation 3 can be divided into two formulations for repeated consumption $R_{t}^{ur}$ and for exploration $R_{t}^{ue}$ recommendations, respectively.
@@
Note that, how to present $R_{t}^{ur}$ and $R_{t}^{ue}$ to users, e.g., as two separate rankings or as a merged ranking of items from both recommendations, is orthogonal to our discussion here.
The partitioning of candidate items into $I^{u}$ and $I^{\bar{u}}$ may again seem trivial (the subscript $t$ is omitted for clarity), but it leads to (i) significant differences in the recommender’s search space, since $|I^{u}|\ll|I^{\bar{u}}|$ and $|I^{\bar{u}}|\approx|I|$, and (ii) consequently, notable impacts on evaluation. Because the recommendation space is much smaller and largely composed of earlier preferred items, making good recommendations among repeated items is far easier than selecting from a large pool of unfamiliar ones. For instance, in food delivery recommendation, the proposed model in (Li et al., 2024) achieved NDCG scores ranging from 0.59 to 0.64 for repeat consumption, while for exploration they ranged only from 0.09 to 0.17 across datasets from three cities. Similar performance gaps were observed for all evaluated baselines (Li et al., 2024). The recommendation accuracy could be easily dominated by the repeated items if repeated consumption is common in the task setting, like food ordering and groceries shopping. Hence, the separation of evaluations of repeated consumption and exploration better reflects the model’s accuracy.
The repeated consumption and exploration is just one example of candidate item selection. The candidate items suitable for recommendation can be directly specified by users e.g., by time, location, or any other item attributes. The key message here is that the set of eligible items to be recommended is determined before running the recommender system.
3.3. Task Formulation with Constraints
We can now define a more general formulation by introducing a selection condition on the candidate items, denoted as $s(I_{\leq t})$. Specifically, $s(\cdot)$ represents selection criteria derived from a user’s interaction history like repeated consumption, or be specified by the user through temporal, spatial, or other attribute-based constraints. By including $s(I_{\leq t})$ as part of the task input, the candidate set of items is determined prior to model ranking. In fact, candidate generation (also known as recall) constitutes the first stage of the multi-stage recommender system architecture widely adopted by today’s largest online platforms, preceding the ranking and re-ranking stages (Liu et al., 2022).